OpenStudy (anonymous):

Part 1: Solve |2x-9|+4 <= 7. You must show all work. Use >= and <= to represent the inequalities. Part 2: Use complete sentences to describe the graph in words. <= (lesser than or equal to) => (greater than or equal to)

OpenStudy (anonymous):

>= (greater than or equal to)*

OpenStudy (anonymous):

Anyone?

OpenStudy (cwrw238):

|2x-9| <= 3 either 2x - 9 <= 3 2x <= 12 or 2x - 9 >= -3 2x >= 6

OpenStudy (anonymous):

So x >= 3

OpenStudy (anonymous):

or x<=6

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

And then how would I descrive the graph?

OpenStudy (cwrw238):

the graph will be a straight line with a clear or filled circle on 3 and 6 and shaded in between filled or clear circles?

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

I don't rmember, I think it's closed circle

OpenStudy (cwrw238):

>= and <= so they will be filled